OpenStudy (anonymous):

The perimeter of rectangle ABCD is 72. The ratio of the lengths of the sides is 1:2. What are the lengths of the sides?

OpenStudy (anonymous):

 So 72 = 2(x) + 2(2x) 72 = 2x + 4x 72 = 6x x = 12 So 12 and 24.